OpenStudy (anonymous):

transform y_1=-4sin(t/2)+3cos(t/2) into y_2=Asin(Bt+C) using y=Msin(Bt)+Ncos(Bt)= sqrt{N^2+M^2}sin(Bt+C) we want to choose C such that |C| is a minimum. so i have so far y_2=5sin(t/2+C). Does C need to be zero?

OpenStudy (anonymous):

\[y=Msin(Bt)+Ncos(Bt)= \sqrt{N^2+M^2}\sin(Bt+C) \]

OpenStudy (anonymous):

compute C to three decimal places so i guess C shouldn't be zero

OpenStudy (anonymous):

and i have figured out that cos(C)=M/sqrt{M^2+N^2} and sin(C)=N/sqrt{M^2+N^2}

OpenStudy (anonymous):

so we need cot(C)=-4/3 so this means C =????? C=-36.870 or -36.87

OpenStudy (anonymous):

that was in degrees i think i will write answer in radians so i think the answer is y=sqrt{M^2+N^2}sin(t/2-.644) what do you guys think? \[y=\sqrt{M^2+N^2}\sin(t/2-.644) \]

OpenStudy (anonymous):

oops M=-4 and N=3 so we have y=5sin(t/2-.644)